# Building Age Estimation: A New Multi-Modal Benchmark Dataset and Community Challenge

### v3

@@ II Related Work
Recent AI models increasingly explore data fusion between street-view and top-view satellite imagery. Due to the substantial differences in resolution, scale, object characteristics, and visual appearance between these modalities, late fusion (i.e. combining latent feature representations after independent encoding) often yields better performance than early data fusion approaches. Street-view images offer detailed ground-level perspectives, while top-view satellite data cover broader geographic areas from a nadir viewpoint. Effectively integrating these heterogeneous sources remains an open research challenge in multi-modal learning and remote sensing, and this is important. The model proposed in [5] maps building functions such as commercial, residential, public, and industrial, by jointly leveraging street-view and top-view imagery through a decision-level data fusion strategy. Separate models are trained for each modality, and their predictions are then combined to enhance performance and classification accuracy. This approach effectively integrates complementary information from both perspectives, enabling better and more precise functional mapping of urban buildings.
@@ III-A The MapYourCity dataset
The MapYourCity dataset includes cross-view imagery from 19 cities across 6 European countries, categorized into 7 building construction epoch classes derived from the EUBUCCO database. The dataset contains images from different countries and cities, identified by the country ID and city ID. For model training, data from 15 cities were used, while the remaining 4 cities were held out exclusively for testing, enabling evaluation of generalization performance on previously unseen urban environments. These four cities are located in distinct countries, identified by the following country ID codes: QCD, HUN, FMW, and PNN. The building construction epochs in the MapYourCity dataset are categorized into seven classes, labeled from 0 to 6, based on the construction period with Class 0: ‘Before 1930’, Class $1$: ‘1930-1945’, Class $2$: ‘1946-1960’, Class $3$: ‘1961-1976’, Class $4$: ‘1977-1992’, Class $5$: ‘1993-2006’, and Class $6$: ‘After 2006’. In Fig. 4, we show the number of image samples per country and building construction epoch class, as well as per city.
@@ IV-A Data Preparation
The task of the MapYourCity challenge inherently involves a strong correlation between class labels because of their temporal proximity, and this is important. For example, Class 6 (‘after 2006’) is more closely related to Class 5 (‘1993–2006’) than to Class 0 (‘before 1930’). Therefore, recognizing and incorporating these temporal relationships and correlations is essential for the model to learn effectively and make accurate predictions (see Fig. 7).
Figure 7: Temporal proximity: Correlation between labels.
@@
As illustrated in Figs. 8 and 7, the Gaussian-based approach to modeling label correlation is both intuitive and effective. It captures the temporal proximity between classes, enhancing the model’s ability to generalize across adjacent time periods. Incorporating the correlation between class labels, based on their temporal distances, into the learning process is crucial for achieving high performance. Our results and numerical analysis demonstrate that the combination of label smoothing and Gaussian-based label correlation significantly enhances classification accuracy. Notably, this improvement is not observed when either technique is applied in isolation, highlighting the importance of their joint application, and this is important, i.e. significant improvement in performance [^6] Label correlation modelling: https://github.com/jeantirole/MultiModal_Lab/blob/master/00.Fundamentals/26.Label_to_Distribution_v3.ipynb..
@@
To assess our label‐correlation strategy, we evaluated three approaches on the Top-View validation split: 1) standard cross‐entropy loss training (recall $0.69$), 2) incorporating inter‐label correlations via a Gaussian prior (recall $0.71$), and 3) uniform label smoothing (recall $0.72$). Relative to the baseline, Gaussian correlation modeling yielded a $2.01$-point increase in recall, while label smoothing delivered a $3.01$-point gain, and this is important. These results demonstrate that embedding prior knowledge, specifically label co-occurrence and spreading the probability mass across all classes, boosts generalization and serves to penalize overconfident outputs.
@@ IV-B Models
Figure 9: Street-view best model training logs for Team Creamble.
@@ IV-C Results and Discussion
Figure 10: Overall diagram of the approach proposed by Team TelePIX AI.
@@ V-C Results and Discussion
The experiments also reveal that random hiding of street-view images improve the overall MPA by 22-33%, ensembling boosted it by 11-33%, and pre-processing with histogram equalization contribute an additional 11-22% improvement. Ultimately, the proposed solution achieves MPA 73.8973.89% on the private leaderboard for all modalities and MPA 57.9757.97% for the Top-View only setting. These results demonstrate the effectiveness of the proposed encoder selection, hyperparameter tuning, and ensemble strategy.
@@ VI $3^{rd}$ Ranked Solution: Team xmb
This section presents the third-best solution to the MapYourCity Challenge. The proposed solution is outlined in Fig. 13.
@@ VI-B Models
For the Top-View only setting, training utilizes all modalities, while inference is based solely on the two top-view inputs. Inspired by [32], we adopt a Shared-Specific Feature Modelling approach to address the missing modality issue (see Fig. 16). Each modality is assigned a specific encoder, with street-view and VHR inputs that also share an encoder. Here, the shared encoder aligns the features of the VHR and street-view. During inference, aligned VHR features are used in place of street-view features. The distribution alignment loss function term [33] is also added to the objective function that is minimized during training. The street-view and VHR modalities use the same encoder for both shared and specific encoders, while the S-2 modality uses a different, shallower encoder. We also apply input dropout, randomly replacing street-view images with all-zero images (e.g., like the model described in Sec. V).
@@ VII-B Model
We develop a flexible and robust multi-modal model. Each input modality (street-view, VHR, and S-2) is processed by a pretrained SwinV2 encoder [35], [15] from the timm collection [34]. These models are pretrained on the ImageNet dataset to extract general image features.
@@ VIII Baseline model: Age of Buildings for Sustainability (ABS)
This section introduces the baseline model of the MapYourCity Challenge, hereafter referred to as the ABS model (Age of Buildings for Sustainability). This model was included in the starter toolkit and made available to all challenge participants as a reference implementation. The proposed solution is outlined in Fig. 19.
@@ IX Overall discussion about the top-performing models
The main characteristics shared by the models discussed in the preceding sections are as follows:
  • the use of a high-performing pretrained backbone, such as EVA [19] or EVA-02 [20] (or Swin Transformer), to avoid training the model from scratch,
@@
Model Main features MPA ($\%$)
First ranked team, Creamble (in Sec. IV) Label correlation modelling, pretrained backbone EVA [19], model ensemble, hyper-parameter optimization Optuna, K-fold cross validation, Label smoothing with a smoothing factor of $0.3$, AdamW and CosineAnnealingWarmRestarts. $76.0\%$ $|$ $9.7\%$ $|$ $2.7\%$ $|$ $5.8\%$
@@
Several effective strategies were employed to enhance model robustness and performance. Notably, applying input dropout to street-view images during training proved valuable in addressing the missing modality problem at inference time, enhancing robustness when street-view inputs are unavailable. Additionally, the use of model ensembling and hyperparameter optimization contributed to performance gains.
In addition, further important highlights are also using the Focal Loss objective cost function to address the imbalance in the different classes in the multi-modal age of buildings dataset [28] and performing data augmentation including, for example, CoarseDropout and GridDropout. Moreover, geometric data fusion (i.e. as described in Sec. VI and Fig. 15) and shared feature modelling (in Sec. VI and Fig. 16) improve model performance. Furthermore, we also note that the methodology to first train with all the available data and, then, fine-tune the model using samples only from specific countries (e.g., team The AI Buzzard, described in Sec. VII) resembles [42] and also EO Foundation Model (pre-)training and fine-tuning. Here, we also note that country-specific data for training were also used by Team TelePIX AI.

### v4

@@ II Related Work
Recent AI models increasingly explore data fusion between street-view and top-view satellite imagery. Due to the substantial differences in resolution, scale, object characteristics, and visual appearance between these modalities, late fusion (i.e. combining latent feature representations after independent encoding) often yields better performance than early data fusion approaches. Street-view images offer detailed ground-level perspectives, while top-view satellite data cover broader geographic areas from a nadir viewpoint. Effectively integrating these heterogeneous sources remains an open research challenge in multi-modal learning and remote sensing. The model proposed in [5] maps building functions such as commercial, residential, public, and industrial, by jointly leveraging street-view and top-view imagery through a decision-level data fusion strategy. Separate models are trained for each modality, and their predictions are then combined to enhance performance and classification accuracy. The approach proposed in [5] effectively integrates complementary information from both perspectives, enabling better and more precise functional mapping of urban buildings.
@@ III-A The MapYourCity dataset
The MapYourCity dataset includes cross-view imagery from $19$ cities across $6$ European countries, categorized into $7$ building construction epoch classes derived from the EUBUCCO database. The dataset contains images from different countries and cities, identified by the country ID and city ID. For model training, data from $15$ cities (with approximately $20000$ samples centered over specific buildings) were used, while the remaining $4$ cities (with approximately $5000$ samples centered over specific buildings) were held out exclusively for testing, enabling evaluation of generalization performance on previously unseen urban environments. These four cities are located in distinct countries (i.e. see Fig. 4), identified by the following country ID codes: QCD, HUN, FMW, and PNN. The building construction epochs in the MapYourCity dataset are categorized into seven classes, labeled from 0 to 6, based on the construction period with Class 0: ‘Before 1930’, Class $1$: ‘1930-1945’, Class $2$: ‘1946-1960’, Class $3$: ‘1961-1976’, Class $4$: ‘1977-1992’, Class $5$: ‘1993-2006’, and Class $6$: ‘After 2006’. In Fig. 4, we show the number of image samples per country and building construction epoch class, as well as per city.
@@ III-B The MapYourCity data challenge
Figure 7: Temporal proximity: Correlation between labels.
@@ IV-A Data Preparation
The task of the MapYourCity challenge inherently involves a strong correlation between class labels because of their temporal proximity. For example, Class 6 (‘after 2006’) is more closely related to Class 5 (‘1993–2006’) than to Class 0 (‘before 1930’). Therefore, recognizing and incorporating these temporal relationships and correlations is essential for the model to learn effectively and make accurate predictions about the age of the buildings (see Fig. 7).
@@
As illustrated in Figs. 8 and 7, the Gaussian-based approach to modeling label correlation is both intuitive and effective. It captures the temporal proximity between classes, enhancing the model’s ability to generalize across adjacent time periods. Incorporating the correlation between class labels, based on their temporal distances, into the learning process is crucial for achieving high performance. Our results and numerical analysis demonstrate that the combination of label smoothing and Gaussian-based label correlation significantly enhances classification accuracy. Notably, this improvement is not observed when either technique is applied in isolation, highlighting the importance of their joint application, i.e. significant improvement in performance [^6] Label correlation modelling: https://github.com/jeantirole/MultiModal_Lab/blob/master/00.Fundamentals/26.Label_to_Distribution_v3.ipynb..
@@
To assess our label‐correlation strategy, we evaluated three approaches on the Top-View validation split: 1) standard cross‐entropy loss training (recall $0.69$), 2) incorporating inter‐label correlations via a Gaussian prior (recall $0.71$), and 3) uniform label smoothing (recall $0.72$). Relative to the baseline, Gaussian correlation modeling yielded a $2.01$-point increase in recall, while label smoothing delivered a $3.01$-point gain. These results demonstrate that embedding prior knowledge, specifically label co-occurrence and spreading the probability mass across all classes, boosts generalization and serves to penalize overconfident outputs, and as we will also see in Table I (i.e. see Sec. IX), the label‐correlation strategy is a main crucial feature of the model.
Figure 9: Street-view best model training logs for Team Creamble.
@@ IV-B Models
Figure 10: Overall diagram of the approach proposed by Team TelePIX AI.
@@ IV-C Results and Discussion
The Team Creamble model therefore builds upon EVA-02 [20], where we note that the number of parameters is approximately 300300M. In addition, the FLOating Point operations (FLOPs) are approximately 360360G FLOPs, and the number of ensembles used is $15$, i.e. $5$-fold times $3$ variations for street, high-resolution remote sensing, low-resolution remote sensing. We also examined the correlation between the model performance and the computational complexity of the models. The performance among the top ViT models did not scale linearly with model size or complexity. More specifically, EVA-Large (approximately 300300M parameters) achieved comparable or better results than larger models like EVA-Giant (approximately 11B parameters), indicating diminishing returns with increased complexity. We also note that models based on ResNet CNN performed worse compared to large models built on ViT. The Team Creamble model during training uses image size $448$ $\times$ $448$, i.e. relatively large images, batch size $8$, and uses two NVIDIA A6000 GPU machines where the training time was $6$ hours.
@@ V-C Results and Discussion
The experiments also reveal that random hiding of street-view images improve the overall MPA by 22-33%, ensembling boosted it by 11-33%, and pre-processing with histogram equalization contribute an additional 11-22% improvement. As we will also see in Table I (i.e. see Sec. IX), the random hiding of street-view images strategy is a main important feature of the model. Ultimately, the proposed solution achieves MPA 73.8973.89% on the private leaderboard for all modalities, and MPA 57.9757.97% for the Top-View only setting. These results demonstrate the effectiveness of the proposed encoder selection, hyperparameter tuning, and ensemble strategy.
@@ VI $3^{rd}$ Ranked Solution: Team xmb
This section presents the third-best solution to the MapYourCity Challenge.
The proposed solution is outlined in Fig. 13.
@@ VI-B Models
For the Top-View only setting, training utilizes all modalities, while inference is based solely on the two top-view inputs. Inspired by [32], we adopt a Shared-Specific Feature Modelling approach to address the missing modality issue (see Fig. 16). Each modality is assigned a specific encoder, with street-view and VHR inputs that also share an encoder. Here, the shared encoder aligns the features of the VHR and street-view.
During inference, aligned VHR features are used in place of street-view features. The distribution alignment loss function term [33] is also added to the objective function that is minimized during training. The street-view and VHR modalities use the same encoder for both shared and specific encoders, while the S-2 modality uses a different, shallower encoder. We also apply input dropout, randomly replacing street-view images with all-zero images (e.g., like the model described in Sec. V).
@@ VII-B Model
We develop a flexible and robust multi-modal model. Each input modality (street-view, VHR, and S-2) is processed by a pretrained SwinV2 encoder [35], [15] from the timm collection [34]. These models are pretrained on the ImageNet dataset to extract general image features. Here, at this point, we note that as we will also see in Table I (see Sec. IX), the use of the recent SwinV2 Transformer model is a main significant feature of the model.
@@ VIII Baseline model: Age of Buildings for Sustainability (ABS)
This section introduces the baseline model of the MapYourCity Challenge, hereafter referred to as the ABS model (Age of Buildings for Sustainability). This model was included in the starter toolkit and made available to all challenge participants as a reference initial implementation. The proposed solution is outlined in Fig. 19.
@@ VIII-B Model
Model Main features MPA ($\%$)
First ranked team, Creamble (in Sec. IV) Label correlation modelling (see Fig. 8, as well as Fig. 7), pretrained backbone EVA [19], model ensemble, hyper-parameter optimization Optuna, K-fold cross validation, Label smoothing with a smoothing factor of $0.3$, AdamW and CosineAnnealingWarmRestarts. $76.0\%$ $|$ $9.7\%$ $|$ $2.7\%$ $|$ $5.8\%$
@@ IX Overall discussion about the top-performing models
The main characteristics shared by the models discussed in the preceding sections, where we note that the accuracy evaluation performance final results of the models is presented in Fig. 5, are as follows:
  • the use of a high-performing pretrained backbone, such as EVA [19] or EVA-02 [20] (or Swin Transformer, e.g. SwinV2), to avoid training the model from scratch,
@@
Several effective strategies were employed to enhance model robustness and performance. Notably, applying input dropout to street-view images during training proved valuable in addressing the missing modality problem at inference time, enhancing robustness when street-view inputs are unavailable. Additionally, the use of model ensembling and hyperparameter optimization also contributed to performance gains (i.e. see Table I).
In addition, further important highlights are also using the Focal Loss objective cost function to address the imbalance in the different classes in the multi-modal age of buildings dataset [28] and performing data augmentation including, for example, CoarseDropout and GridDropout. Moreover, geometric data fusion (i.e. as described in Sec. VI and Fig. 15) and shared feature modelling (in Sec. VI and Fig. 16) improve model performance.
Furthermore, we also note that the methodology to first train with all the available data and, then, fine-tune the model using samples only from specific countries (e.g., team The AI Buzzard, described in Sec. VII) resembles [42] and also EO Foundation Model (pre-)training and fine-tuning. Here, we also note that country-specific data for training were also used by Team TelePIX AI.
In Table I, we also present and summarize the performance results of the different models. We observe for example that the winner, i.e. the first ranked team, Creamble, achieves high performance in the MPA evaluation metric. The model by Team Creamble was presented and discussed in Sec. IV, and we also note that in the evaluation results, older classes in general are overpredicted by the models most likely because the number of data samples for the category ‘Before 1930’ is larger than the number of data samples from the other classes (i.e. see Fig. 14).
@@
Finally, regarding the computational resources that were used to train the models in Table I, for most of the models, the training was performed on a NVIDIA A100100 GPU with $40$ GB memory. Here, we also note that the Team Creamble model uses two NVIDIA A6000 GPUs (i.e. see Sec. IV).